### v1

# Explanation Regularisation through the Lens of Attributions

Abstract
Explanation regularisation (ER) has been introduced as a way to guide models to make their predictions in a manner more akin to humans, i.e., making their attributions ‘plausible’. This is achieved by introducing an auxiliary explanation loss, that measures how well the output of an input attribution technique for the model agrees with relevant human-annotated rationales. One positive outcome of using ER appears to be improved performance in out-of-domain (OOD) settings, presumably due to an increased reliance on ‘plausible’ tokens. However, previous work has under-explored the impact of the ER objective on model attributions, in particular when obtained with techniques other than the one used to train ER. In this work, we contribute a study of ER’s effectiveness at informing classification decisions on plausible tokens, and the relationship between increased plausibility and robustness to OOD conditions. Through a series of analyses, we find that the connection between ER and the ability of a classifier to rely on plausible features has been overstated and that a stronger reliance on plausible tokens does not seem to be the cause for any perceived OOD improvements. [^1] Source code available at https://github.com/PedroMLF/ER_through_the_lens_of_attributions.
Explanation Regularisation through the Lens of Attributions
Pedro Ferreira 1  Wilker Aziz 1  Ivan Titov 1,2 1 University of Amsterdam  2 University of Edinburgh { p.m.ferreira, w.aziz }@uva.nlititov@inf.ed.ac.uk
1 Introduction
Figure 1: ER (top-left) minimises a classification and an explanation loss. The latter uses an input attribution technique to obtain a machine rationale for the prediction, and penalises differences between that and the human rationale. Guiding the model to attribute predictions to plausible tokens is expected to help the classifier after training (top-right), when no human rationale is available, and the classifier might be exposed to OOD inputs. Bottom: as any attribution technique is only a proxy to faithful input attributions, there can be a mismatch between predicting the human rationale (through the guided attribution technique) and effectively informing the classifier’s prediction by human rationale tokens.
In explanation regularisation (ER; Ross et al., 2017; Ghaeini et al., 2019; Liu and Avci, 2019; Ismail et al., 2021; Joshi et al., 2022; Pruthi et al., 2022; Stacey et al., 2022, i.a.), a text classifier is encouraged to inform its predictions by tokens identified as part of a human rationale for the label. [^2] In this setting, a human rationale is the subset of input tokens that the annotator perceives as relevant to justify the label. This is achieved with an auxiliary ‘explanation loss’ that penalises differences between the output of a guided attribution technique (e.g., a relevance map based on top-layer attention) and the annotated human rationale (see Figure 1). Compared to their baseline counterparts trained without rationales, these ER models have been reported to improve classification in several out-of-domain (OOD) conditions, e.g., unseen test data (Joshi et al., 2022; Stacey et al., 2022; Madani and Minervini, 2023). This improved robustness is ascribed to increased reliance of ER models on plausible tokens (i.e., those in the human rationale for the label).
While ER is assumed to encourage classifiers to attribute their decisions more plausibly, and this increased plausibility is assumed to explain the observed OOD improvements, these have not been carefully verified in previous work. In particular, existing analyses have focused on the impact of explanation regularisation on guided attributions (i.e., those provided by the technique used in training), finding that ER models are better at aligning those with human rationales (Stacey et al., 2022; Madani and Minervini, 2023). However, through the explanation loss, guided attributions are explicitly supervised to resemble human rationales. Hence, we argue that this observation is an expected consequence of minimising the explanation loss, and not evidence that the classifier attributes its decisions more plausibly. Instead, we need to entertain the possibility of a faithfulness gap: the output of an attribution technique might not necessarily be aligned with the decision process of the model. Furthermore, ER can exacerbate this gap, making the guided attribution technique even less faithful.
In this work, we analyse (i) ER’s ability to make classifiers rely on plausible tokens; and (ii) the relationship between increased plausibility and robustness to OOD conditions. We start by studying in an in-domain setting how ER impacts input attributions using a range of post-hoc attribution techniques (§6.2). Here we find that, while ER can find models whose guided attributions effectively predict human rationales, there is little to no evidence of that being true under other (evaluation) post-hoc attribution techniques. This makes us doubt that ER substantially affects the model’s classification decision process. This finding persists even after making use of a potentially more faithful guided attribution technique (§6.3). Only by strongly constraining the model to achieve a low explanation loss in training, do we observe increased reliance on plausible features (§6.4), but that happens at the expense of OOD improvements. We follow this result by studying the trade-off between the classification and explanation losses (§6.5). Finally, we look into whether in-domain classification performance and plausibility of attributions available for model selection can be predictive of OOD classification improvements (§6.6). We observe that OOD improvements are difficult to predict, and that ER’s impact on the plausibility of input attributions of OOD examples is more limited than in-domain.
To summarise, we find that the connection between ER and the ability of a classifier to rely on plausible tokens has been overstated. While there is likely some effect connecting OOD classification improvements to classifiers that are better predictors of plausible tokens, the evidence we collect points to this effect not being a stronger reliance on plausible tokens.
2 Related Work
Learning from explanations.
Human explanations, e.g., rationales, have been used to improve text classifiers (Hartmann and Sonntag, 2022; Hase and Bansal, 2022). Examples include select-predict ‘pipelines’, where an explanation output by a first module serves as input to a classifier (Camburu et al., 2018), and multi-task settings, where a classifier and an explanation predictor model are jointly trained (Chan et al., 2022). The latter can also be framed as a multi-objective optimization problem, as in Resck et al. (2024), a recent work on the trade-off between plausibility and in-domain classification performance in Pareto-optimal models.  In contrast, we are interested in explanation regularisation, where a model is trained such that its attributions, according to a differentiable attribution technique, match existing rationales (Ross et al., 2017; Ghaeini et al., 2019; Liu and Avci, 2019; Rieger et al., 2020; Mathew et al., 2021; Pruthi et al., 2022; Stacey et al., 2022, i.a.). Most work on ER focuses on in-domain data, however, there is evidence that ER is capable of improving robustness to OOD data, for both text Rieger et al. (2020); Joshi et al. (2022); Stacey et al. (2022), and image classification Chefer et al. (2022); Rao et al. (2023). We focus on this setup and further explore how ER impacts OOD attributions and performance.
Attribution techniques.
In vector-based techniques, model components are directly used to obtain input attributions. The simplest approach corresponds to using attention weights (Clark et al., 2019, i.a.), despite conflicting evidence on its usefulness for this purpose (Jain and Wallace, 2019; Serrano and Smith, 2019; Wiegreffe and Pinter, 2019). Later works (Kobayashi et al., 2020, 2021; Ferrando et al., 2022; Modarressi et al., 2022) incorporate information about the magnitude of input vectors and the influence of other parts of the Transformer layer (Vaswani et al., 2017). These are examples of local techniques, where attributions are based on the dynamics of a single layer.
It is also possible to obtain a global analysis of the model, by including the dynamics of its multiple layers. One example is attention-rollout (Abnar and Zuidema, 2020), which recursively aggregates attention weights across layers. Other examples of global vector-based techniques are GlobEnc (Modarressi et al., 2022), ALTI (Ferrando et al., 2022), and DecompX (Modarressi et al., 2023). Global attributions can also be obtained from gradient-based approaches (Kindermans et al., 2016; Shrikumar et al., 2017; Sundararajan et al., 2017, i.a.), or a mix of attention and gradients (Chefer et al., 2021; Qiang et al., 2022).
3 Explanation Regularisation
In ER (Joshi et al., 2022; Stacey et al., 2022; Pruthi et al., 2022, i.a.), we assume the availability of training data $\mathcal{D}=\{(x_{i},y_{i},e_{i})_{i=1}^{N}\}$, for each data point, $x_{i}$ is the input text, $y_{i}$ is the target label, and $e_{i}$ is a human rationale (specifically, a subset of tokens of the input regarded as a justification for the label). Human rationales can be obtained from annotators (Socher et al., 2013; Camburu et al., 2018; Rajani et al., 2019; Mathew et al., 2021) or through some form of heuristic to select relevant tokens (Liu and Avci, 2019; Rieger et al., 2020), and are only required during training.
An ER classifier minimises a joint loss
$\mathcal{L}_{\mathrm{cls}}(\theta|\mathcal{D})+\lambda~{}\mathcal{L}_{\mathrm{expl}}(\theta|\mathcal{D}),$ (1)
where $\mathcal{L}_{\mathrm{cls}}$ is the regular classification loss and $\mathcal{L}_{\mathrm{expl}}$ is the explanation loss, with a hyperparameter $\lambda\in\mathbb{R}_{\geq 0}$ controlling the contribution of the latter. The expectation is that $\mathcal{L}_{\mathrm{expl}}$ encourages the model to attribute classification decisions more similarly to human rationales, and this, in turn, leads to models that achieve better classification performance, particularly in out-of-domain conditions.
The first term in Eq. (1) is the cross-entropy loss
$\mathcal{L}_{\mathrm{cls}}(\theta|\mathcal{D})=\mathbb{E}_{(x,y)\sim\mathcal{D}}[\mathcal{L}_{\mathrm{ce}}(C_{\theta}(x),y)]~{},$ (2)
where $C_{\theta}$ is a trainable classifier. [^3] The typical $C_{\theta}$ employs a pre-trained text encoder and projects the top-layer representation of the [CLS] token to a probability distribution over all classes. The second term employs an attribution technique in order to obtain a relevance map $E(C_{\theta},x)$ for the classifier’s output. ER takes this as a proxy to faithful attributions, and penalises the classifier by this map’s deviation from the human rationale, as assessed by a function $\mathcal{S}$ such as mean absolute error, mean squared error, or KL divergence (Joshi et al., 2022):
$\mathcal{L}_{\mathrm{expl}}(\theta|\mathcal{D})=\mathbb{E}_{(x,y,e)\sim\mathcal{D}}[\mathcal{S}(E(C_{\theta},x),e)].$ (3)
To avoid confusing the posthoc attribution technique that guides the model during training with other posthoc attribution techniques we may be interest in for analysis, we will refer to the former as the guided attribution technique and the latter as the evaluation attribution techniques. In addition to being differentiable, the guided attribution technique is required to be memory- and time-efficient, since it is used for training. Examples of such techniques used in ER include gradient-based (Ross et al., 2017; Ghaeini et al., 2019; Chefer et al., 2022), vector-based (Mathew et al., 2021; Pruthi et al., 2022; Stacey et al., 2022; Fernandes et al., 2022), and perturbation-based approaches (Ying et al., 2022). We follow Joshi et al. (2022) and Stacey et al. (2022) and use top-layer attention.
4 Method
We now outline two changes to the ER approach (§3) that provide the necessary level of empirical evidence to study our research questions.
4.1 Global Attributions as Supervision Signal
Intuitively, the guided attribution technique in ER controls the potential outcome of exploiting the additional supervision in human rationale data: if highly faithful to the classifier’s decision process, the model will be constrained to processing the input as if it were masked (i.e., to omit tokens the annotator deemed unessential), otherwise, the model will be able to process the input somewhat independently of what the guided attributions look like (i.e., of how human-rationale-looking they happen to be). A potential side effect of the former is to overly-constrain the classifier, forcing it to disregard portions of the input necessary for (machine) decision making. A potential side effect of the latter is to make no use of the additional supervision (e.g., the model might have enough capacity to minimise $\mathcal{L}_{\text{expl}}$ without any meaningful changes to the features available to the classification head). To study this, we employ a global attribution technique for guidance. For taking more of the model’s components into account when building a relevance map for the prediction, these techniques are potentially more faithful, or, as at least intuition suggests, they may lead to more substantial changes to how the input is processed at lower layers. We experiment with attention rollout (Abnar and Zuidema, 2020) for, on the one hand, it is a direct global counterpart to top-layer attention, and, on the other hand, it is much more memory- and time-efficient than vector-based alternatives (e.g., ALTI and DecompX). In Appendix A, we develop the intuition a bit more by comparing $\nabla_{\theta}\mathcal{L}_{\text{expl}}$ as a function of whether the guided attribution technique is local or global.
4.2 Constrained Optimization Approach
Suppose we are interested in analysing the best classifier we can get while meeting a constraint on how it attributes its decisions to input tokens. This is useful, for example, when studying the connection between reliance on plausible (human-rationale) tokens and OOD performance across different design choices (e.g., guided attribution technique). We can reinterpret the ER objective slightly and achieve that, by casting it as a constrained optimisation problem—minimise classification loss subject to a bound on explanation loss:
$\min_{\theta}$ $\quad\mathcal{L}_{\mathrm{cls}}(\theta|\mathcal{D})$ $\mathrm{s.t.}$ $\quad\mathcal{L}_{\mathrm{expl}}(\theta|\mathcal{D})\leq b\;.$ (4)
For each choice of guided attribution technique, we can obtain a bound $b$ by interpreting $\mathcal{L}_{\text{expl}}$ as the sole objective of optimisation. To select the bound, we first independently minimise $\mathcal{L}_{\mathrm{expl}}$, setting $b$ to $\min_{\theta}~{}\mathcal{L}_{\mathrm{expl}}(\theta|\mathcal{D})$. For Eq. (4), we use Lagrangian relaxation (Boyd and Vandenberghe, 2004). Implementation choices are detailed in Appendix F.
5 Experimental Setting
Data.
We use SST-2 (Socher et al., 2013) as training and in-domain data, following the heuristic algorithm proposed in Carton et al. (2020) to obtain instance-level rationales. For OOD we use: Amazon Reviews (Hou et al., 2024), IMBD (Maas et al., 2011), Movies (Zaidan and Eisner, 2008; DeYoung et al., 2020), and Yelp (Zhang et al., 2015). For more details on data refer to Appendix F.2.
Model
All experiments use HuggingFace Transformers’ (Wolf et al., 2020) BigBird-roberta-base model as the pre-trained contextual embedding encoder model (Zaheer et al., 2020). This choice follows previous works (Chan et al., 2022; Joshi et al., 2022; Madani and Minervini, 2023). For more training details refer to Appendix F.
Scoring Function.
We use Mean Absolute Error (MAE) as the scoring function $\mathcal{S}$ in Equation 3.
Guided Attribution Technique.
We use top-layer attention scores (Att) as our local guided attribution technique, $E_{c}$, as explored before in the literature (Joshi et al., 2022; Stacey et al., 2022; Pruthi et al., 2022). As part of our analysis we also use attention rollout (AttR) as a global guided attribution technique. This technique recursively aggregates the local attention scores into a global representation. In both cases we use the scores for the [CLS] token, averaged across heads.
Target Annotations.
All datasets with rationales are annotated at the instance level, with a binary vector with length equal to the number of tokens, where a value of one corresponds to a token that is highlighted as relevant. In order to obtain a target for the scoring function $\mathcal{S}$, the instance-level vector is normalized to sum to one by dividing each element by the total number of highlighted tokens in the instance (Joshi et al., 2022; Stacey et al., 2022).
ID OOD
Dev Test Movies Yelp IMDB Amazon-M-TV
Baseline 92.95 $\pm$ 0.57 94.36 $\pm$ 0.48 91.01 $\pm$ 4.38 94.41 $\pm$ 0.99 91.20 $\pm$ 1.63 83.89 $\pm$ 0.89
Joint6.2 and §6.3)
Attention 93.56 $\pm$ 0.50 94.33 $\pm$ 0.72 93.10 $\pm$ 2.44 95.09 $\pm$ 0.44 91.87 $\pm$ 1.10 84.10 $\pm$ 0.56
+ Rollout 93.19 $\pm$ 0.68 94.09 $\pm$ 0.82 92.72 $\pm$ 2.71 94.77 $\pm$ 0.66 91.88 $\pm$ 0.99 84.15 $\pm$ 0.47
Constrained6.4)
Attention 93.41 $\pm$ 0.57 94.45 $\pm$ 0.49 90.59 $\pm$ 4.80 94.52 $\pm$ 1.30 90.65 $\pm$ 2.17 83.38 $\pm$ 1.11
+ Rollout 90.78 $\pm$ 0.90 91.11 $\pm$ 0.77 89.56 $\pm$ 2.67 92.40 $\pm$ 0.78 89.19 $\pm$ 0.85 81.44 $\pm$ 0.81
Table 1: F1-Macro scores ($\uparrow$) and standard deviation for in-domain (ID) and out-of-domain (OOD) data. Attention corresponds to a model that uses attention as the guided attribution technique, and +Rollout to a model that uses attention-rollout. Results are averages of 15 seeds.
Figure 2: F1-Macro scores ($\uparrow$) for both in-domain and out-of-domain data. ER-A is an ER model that uses attention as the guided attribution technique, and ER-R is an ER model that uses attention-rollout. The C prefix indicates a constrained model. Results correspond to 15 seeds.
5.1 Attribution Techniques
We evaluate the plausibility of attributions using four post-hoc attribution techniques to compute token-level global attributions: InputXGradient (IxG; Shrikumar et al., 2017), a gradient-based method that obtains attributions by multiplying inputs by their respective gradients; ALTI Ferrando et al. (2022), a vector-based technique that measures token-to-token interactions at a layer level,and that linearly combines these interactions to provide global attribution scores; and DecompX (DX; Modarressi et al., 2023), a global vector-based technique based on the propagation of decomposed token representations through the model (we also use a variant, DX-C, that includes the classification head). For IxG we use the Captum package (Kokhlikyan et al., 2020). For DX, DX-C [^4] https://github.com/mohsenfayyaz/DecompX and ALTI [^5] https://github.com/mt-upc/transformer-contributions we adapt the available code.
5.2 Plausibility Metrics
To measure how well the post-hoc input attributions reproduce the human annotated rationales, we employ three metrics introduced by Fomicheva et al. (2021), and further described in Appendix B: AUC Score, Average Precision; and Recall@k.
6 Results
6.1 Classification Task Generalisation
We start by reporting how ER+Att classification performance generalises to OOD data. Table 1 shows the average F1-Macro scores for both in- and OOD data. Similarly to Joshi et al. (2022) and Stacey et al. (2022), we observe a positive impact in OOD average classification performance when training with ER. However, if we consider standard deviation, and visualise the distribution of results for all model seeds (Figure 2), it is noticeable how the effect on the reported average improvements are exacerbated by certain model versions that perform poorly. In fact, this seems to indicate that one of the merits of ER is to converge less often to models that generalise poorly to OOD conditions.
6.2 ER Impact on Model Attributions
Existing work on explanation regularisation has shown evidence of improved alignment between the output of the guided attribution technique and human annotated rationales (Stacey et al., 2022; Madani and Minervini, 2023). However, this analysis is limited to the guided attribution technique (i.e., the one used during training). We argue that such improvements are to be expected, provided that the model converges to a solution that minimizes the explanation loss. However, attributions revealed by an attribution technique are not necessarily aligned with the model’s decision process—there exists a faithfulness gap for input attribution methods. The increased alignment between the guided attention technique and the human rationale may have been achieved by broadening the gap rather than making the ER model more reliant on plausible tokens. In order to test this hypothesis, we compute plausibility metrics (Section 5.2) for attributions with a series of techniques described in Section 5.1. If ER genuinely improves model reliance on plausible features then we would expect to observe changes in attributions for other techniques, rather than only for the one used in training. We show average scores for AUC in Table 2.
As expected, ER+Att is able to strongly impact the average AUC plausibility of the guided attribution technique (Att), by increasing its value from 35.6 (in the baseline) to 78.6. However, for the remaining suite of attribution techniques, the impact is small or non-existent (e.g., 50.6 vs. 51.5 for IxG). Thus, we can indeed conclude that the improvement in plausibility for attribution obtained with Att was achieved largely by broadening the faithfulness gap rather than by making the ER model more reliant on plausible features. [^6] We find similar evidence using average precision and recall@k plausibility metrics (Appendix Table 8).
SST-Dev
Att
AttR
IxG
ALTI
DX-C
DX
Baseline 35.6 27.9 50.7 59.0 58.5 57.5
Joint ER6.2 and §6.3)
w/ Att 78.6 27.9 50.6 64.5 58.4 57.8
w/ AttR 37.0 28.7 50.8 60.2 58.6 57.4
Constrained ER6.4)
w/ Att 87.6 28.2 51.5 68.3 57.6 56.8
w/ AttR 63.2 89.1 67.6 87.9 76.9 81.3
$\mathcal{L}=\mathcal{L}_{\mathrm{expl}}$
$\mathcal{L}_{\text{expl}}$ (A) 89.0 29.6 53.2 78.2 52.0 53.0
$\mathcal{L}_{\text{expl}}$ (R) 81.1 89.4 72.9 88.2 77.1 83.3
Table 2: Average AUC plausibility scores ($\uparrow$). $\mathcal{L}_{\text{expl}}$ (A) and $\mathcal{L}_{\text{expl}}$ (R) are trained with $\mathcal{L}=\mathcal{L}_{\mathrm{expl}}$, for Att and AttR, respectively. Highlighted values correspond to the guided attribution technique used for ER.
6.3 ER with Global Attribution Technique
We could hypothesise that the failure to enhance the genuine reliance of the method on plausible features is due to the use of attention in ER. It is evident that top-layer attention can align with human rationales without altering the underlying function, for example, by rescaling its input vectors to compensate for the change in attention or by shifting ‘implausible’ computations to a lower layer, thus, hiding them from the regulariser. Therefore, here, we test whether the same behavior is observed with a more reliable global attribution technique. We choose attention-rollout (AttR; Abnar and Zuidema, 2020), a computationally efficient way of obtaining differentiable global vector-based attributions (see Section 4.1).
In terms of classification performance, ER+ AttR performs similarly to ER+Att (Table 1). Also, despite its potential to lead ER models to rely on plausible features, we find once again that other attributions are not affected (Table 2). This seems to indicate that a joint loss setup in ER does not result in models that lead to more plausible attributions overall. It is also noticeable how, contrarily to what we have seen with ER+Att in Section 6.2, ER+AttR almost did not impact the respective guided attribution technique (28.7 vs 27.9). To understand why this might have happened, we train ER models exclusively with the explanation loss, $\mathcal{L}=\mathcal{L}_{\mathrm{expl}}$, and obtain ‘optimal’ loss values. We then compare these with the values obtained for the joint ER models. As observed in Table 3, there is a gap to the ‘optimal’ explanation loss value for both approaches, more so for AttR, extending to AUC—for $\mathcal{L}_{\text{expl}}$ (A) we obtain a AUC value of 89 (vs the 78.6 obtained with joint ER+Att), and for $\mathcal{L}_{\text{expl}}$ (R) we obtain a value of 89.4 (vs the 28.7 obtained with joint ER+AttR). This shows that ER models optimised with a joint objective can converge to a solution where the model effectively ignores the auxiliary explanation loss component.
$E_{c}$ Objective $\mathcal{L}^{(\text{dev})}_{\mathrm{expl}}$  ($\downarrow$) AUC${}^{(\text{dev})}$ ($\uparrow$)
Att $\mathcal{L}_{\text{expl}}$ 0.030 89.0
(J) ER 0.040 78.6
(C) ER 0.034 87.6
AttR $\mathcal{L}_{\text{expl}}$ 0.032 89.4
(J) ER 0.064 28.7
(C) ER 0.033 89.1
Table 3: SST-Dev explanation loss and AUC values for models guided with attention (Att) or attention-rollout (AttR), as a function of their training objective (explanation loss alone, vs. (J)oint ER and (C)onstrained ER).
6.4 Constrained Optimization Approach
We have observed how the joint objective might result in ER models that ‘under-optimise’ the explanation loss. Yet, the results shown in Table 2 for the models trained exclusively with the explanation loss, $\mathcal{L}=\mathcal{L}_{\mathrm{expl}}$, show an interesting observation: attention-rollout as the guided attribution technique leads to a model that more strongly relies on plausible tokens, as observed by the positive impact across all reported attribution techniques. In order to bridge the gap between these two observations, we make use of constrained optimisation (Section 4.2). The goal is that, by constraining ER to converge to solutions that are close to the ‘optimal’ explanation loss, while still optimising for classification performance, we can better evaluate the impact of using models whose classification decisions rely more on plausible tokens.
As observed in Table 2, while constrained ER+Att still mostly fails to impact the evaluation attribution techniques, constrained ER+AttR impacts the plausibility of all assessed techniques. However, despite the positive impact of constrained ER+AttR on the plausibility of attributions, Table 1 indicates a decline in classification performance. In summary, the only ER strategy we tested that results in a model relying more strongly on plausible tokens decreases OOD classification robustness. This suggests a disconnect between (over-)reliance on plausible features and OOD robustness.
6.5 Classification/Explanation Loss Trade-off
Figure 3: Cross-Entropy vs Explanation Loss for SST-Dev. The vertical lines show the explanation-loss bound. Each point corresponds to the average of 5 runs, and the error bars to the standard deviation for each loss. $\lambda$ ranges from 0 (No-ER) to 100.
We have observed how constrained optimisation leads to models that are able to retain some of the classification task performance, but do not improve the plausibility of input attributions (Constrained ER+Att), or models that do improve the plausibility of attributions, but at the expense of the classification task performance (Constrained ER+AttR). In order to better understand how the choice of ER attribution technique impacts the trade-off between the two losses we are jointly optimising, we study this scenario with increasing values of $\lambda$.
We observe two overall trends for SST-Dev data (Figure 3). Given a high enough $\lambda$, the explanation loss converges to the ‘optimal’ empirical value. For ER+Att, smaller values of $\lambda$ are enough to impact the explanation loss, whereas for ER+AttR larger values are required. For both strategies alike, as the explanation loss values decrease, the cross entropy loss degrades (we do verify that all models are still learning to solve the classification task). However, the ER+Att model seems to be more robust to this trade-off—we observe a smaller drop in CE loss as the explanation loss decreases when compared to ER+AttR. This is aligned with our previous observation: improving the explanation loss (i.e. the model is better at predicting the human rationales) negatively impacts the CE loss, decreasing classification task performance, where ER+Att outperforms ER+AttR (Table 1).
The presence and the prominence of this trade-off may seem puzzling. We hypothesise two potential, non-mutually exclusive explanations. First, human rationales may not accurately reflect features useful for models due to noise in annotation or the inability of annotators to identify all genuinely predictive features (Carton et al., 2020). Second, attention regularisation imposes assumptions about how the model should implement the function (e.g., aggregating information in the top layer for ER+Att), which may be sub-optimal for the task. Investigating these hypotheses requires a range of experiments across various tasks and models, which is beyond the scope of this paper. However, we can already conclude that the presence of the trade-off reduces the opportunity to simultaneously improve both OOD performance and reliance on plausible features.
6.6 Predicting OOD Performance
As shown in Figure 2, ER models exhibit a broad spread of OOD classification results, highlighting the importance of identifying the best-performing models, ideally without access to OOD data. With that goal, we investigate whether it is possible to predict model performance, by studying if (i) in-domain classification or plausibility performance predicts OOD classification performance; (ii) OOD plausibility predicts OOD classification performance; and (iii) the auxiliary task transfers OOD.
(a) Predictor: SST-Dev Classification
(b) Predictor: SST-Dev Plausibility AUC w/ DX-C
(c) Predictor: Movies / Yelp-50 Plausibility AUC w/ IxG
Figure 4: Relationship between ID/OOD predictors and OOD classification performance.
From In-Domain Performance.
Figures 4(a) and 4(b) show the association between in-domain (SST-Dev) predictors classification F1-Macro and plausibility AUC (computed with DX-C), and classification F1-Macro for SST-Test (in-domain) and Movies (OOD). In both cases there is no clear correlation between the in-domain metrics and OOD classification scores, with the result in Figure 4(b) further supporting the apparent disconnect between reliance on plausible features and OOD classification performance. We find a similar behavior for other approaches, OOD datasets, plausibility metrics and attribution techniques (Appendix D).
From OOD Plausibility Performance.
Given that we do not find in-domain classification performance or attribution plausibility to be predictive of OOD performance, we now investigate if there is any correlation at the OOD level. In order to test this, we use the Movies dataset, and Yelp-50, a subset of Yelp, annotated as described in Appendix E. Figure 4(c) shows the association between OOD plausibility AUC (computed with IxG [^7] Movies consists of long inputs, resulting in prohibitive runtime for batch sizes that do not result in OOM errors, on a Nvidia A100, when using DX and ALTI. Thus, we report IxG, as it can be computed for both datasets in the figure.) and classification F1-Macro, for both datasets. Similarly to what we observed in-domain, there is no correlation between OOD plausibility scores and its corresponding classification task performance.
ER Impact on OOD Attributions.
If improved OOD classification performance is to be ascribed to more plausible attributions, we would expect ER to impact plausibility metrics OOD. Results for AUC (Table 4) show that the impact of ER on OOD attributions is still present, but more limited than in-domain (Table 2), hinting at a limited OOD impact of the auxiliary task introduced in ER.
Movies Yelp-50
Att
AttR
IxG
Att
AttR
IxG
Baseline 70.0 49.6 58.1 51.9 37.1 51.5
Joint ER6.2 and §6.3)
w/ Att 74.6 49.1 58.1 69.7 36.4 50.7
w/ AttR 70.1 50.2 58.0 52.5 37.6 51.7
Constrained ER6.4)
w/ Att 68.8 48.1 56.0 76.0 37.4 51.3
w/ AttR 71.1 62.8 59.3 67.2 69.4 58.8
$\mathcal{L}=\mathcal{L}_{\mathrm{expl}}$
$\mathcal{L}_{\text{expl}}$ (A) 61.0 48.2 48.0 70.1 38.5 47.1
$\mathcal{L}_{\text{expl}}$ (R) 59.4 62.5 55.2 62.6 68.6 56.8
Table 4: Average AUC plausibility scores ($\uparrow$). $\mathcal{L}_{\text{expl}}$ (A) and $\mathcal{L}_{\text{expl}}$ (R) are trained with $\mathcal{L}=\mathcal{L}_{\mathrm{expl}}$, for Att and AttR, respectively. Highlighted values correspond to the guided attribution technique used for ER.
7 Conclusion
In this work we take a step towards a better understanding of how explanation regularisation impacts a model beyond its classification performance. In particular, we aim to study both in- and out-of-domain settings, with a focus on how the plausibility of input attributions is affected, and what is the relationship between increased reliance on plausible tokens and robustness to OOD conditions.
We find that ER, unless constrained to meet an ‘optimum’ explanation loss computed with a global attribution technique, does not lead to models that effectively attribute to more plausible tokens. In fact, we observe a disconnect between reliance on plausible tokens and OOD robustness, with (i) OOD classification improvements not being predicted by increased plausibility; and (ii) OOD classification performance degrading when models are constrained to rely more on plausible tokens.
Our findings highlight relevant challenges for ER and suggest future research questions: (i) how to achieve a better balance between constraining the model to attend to plausible tokens and solving the classification task, e.g., by learning which layers and attention heads of the Transformer model to regularise (see (Fernandes et al., 2022)), or (ii) by allowing the model to learn to select which rationales to use during training (see (Arous et al., 2021; Carton et al., 2022)); and (iii) how would input attributions for an ER model be impacted by using more faithful, while still efficient, vector-based global guidance techniques as part of ER design choices (and not only as an analysis tool).
Limitations
In this section we identify some limitations of our work. However, we do note that we focus on the limitations that stem from design choices of our analysis, and not on the limitations we ‘inherit’ and that are part of the base problem, e.g., whether plausibility should be a desired property of explanations (Jacovi and Goldberg, 2021).
Models.
We use a single pre-trained model, BigBird-roberta-base (Zaheer et al., 2020). This choice follows existing work on ER (in particular applied to out-of-domain conditions), and allows evaluating datasets with long examples, but it does mean that we do not assess how ER impacts attributions for different pre-trained models, in particular, with varying number of parameters.
Data.
In terms of data, we identify two independent limitations. The first limitation related with data concerns the limited number of datasets per task that include annotations, making it difficult to analyse the impact of ER on the attributions on out-of-domain settings. Secondly, we use only the task of sentiment analysis to study the disconnect between ER’s robustness to OOD conditions and the plausibility of attributions. In practice, it would be possible to conduct a similar analysis for other datasets that include human annotations. However, the questions we ask are deeply inherent to ER, and it seems unlikely that alternative tasks would be impacted differently.
Attribution Techniques.
Two of the techniques we use, ALTI (Ferrando et al., 2022) and DecompX (Modarressi et al., 2023), cannot be applied to long examples, leading either to prohibitive runtime, or to out-of-memory errors (on a single GPU NVIDIA A100), even with a batch size as low as 2. Thus, we cannot apply them to the OOD movies dataset, and have to limit our manually annotated split of Yelp to a moderate maximum sequence length. Yet, on Yelp, these represent a reasonable portion of the dataset’s length distribution.
Ethical Considerations
This work studies the synergy between models that attribute more plausibly, i.e., more aligned with humans, and improved OOD performance. We use existing datasets, that include human annotations of the relevant tokens that explain why an example is classified with a given label. These annotations might be biased, include mistakes, etc. By training a text classifier to become more aligned with those annotations we might further magnify the biases of the annotation process.
Acknowledgements
This research was done within the Mercury Machine Learning Lab, a collaboration between the University of Amsterdam, TU Delft, and Booking.com. Ivan Titov is supported by the Dutch National Science Foundation (NWO Vici VI.C.212.053). All content represents the opinion of the authors, which is not necessarily shared or endorsed by their respective employers and/or sponsors.
Appendix A Attention-Rollout
Rollout was introduced in Abnar and Zuidema (2020), and makes it possible to obtain global attributions by recursively aggregating vector-based local attributions, such as attention. We can also use it to show why we expect global attributions to be more impactful than the local counterparts when supervising a model with explanations. By defining attention-rollout recursively, $\mathbf{a}_{l}=\rho(\mathbf{h}^{l},\mathbf{h}^{l-1};\mathbf{a}_{l-1})$, where $\mathbf{a}_{l}$ corresponds to attention-rollout weights at layer $l$, and $\mathbf{h}^{l}$ to the hidden states at layer $l$, we can write its gradient. Namely, for the $l$th layer, we get:
$\partialderivative{\mathbf{a}_{l}}{\theta}=$ $\partialderivative{\theta}\rho(\mathbf{h}^{l},\mathbf{h}^{l-1};\mathbf{a}_{l-1})$ (5)
$=$ $\partialderivative{\mathbf{h}^{l}}\rho(\mathbf{h}^{l},\mathbf{h}^{l-1};\mathbf{a}_{l-1})\times\partialderivative{\mathbf{h}^{l}}{\theta}$
$+$ $\partialderivative{\mathbf{h}^{l-1}}\rho(\mathbf{h}^{l},\mathbf{h}^{l-1};\mathbf{a}_{l-1})\times\partialderivative{\mathbf{h}^{l-1}}{\theta}$
$+$ $\partialderivative{\mathbf{a}_{l-1}}\rho(\mathbf{h}^{l},\mathbf{h}^{l-1};\mathbf{a}_{l-1})\times\underbrace{\partialderivative{\mathbf{a}_{l-1}}{\theta}}_{\text{recursion}}.$
By inspecting Equation 5, we can observe how the impact of guiding with a global vector-based technique goes beyond that of a local technique – there we have $\mathbf{a}_{l}=\rho(\mathbf{h}^{l},\mathbf{h}^{l-1})$, meaning that the last recursion term of the gradient that propagates to the layers below would not be part of the computation. This difference seems to indicate that using a global attribution technique, such as attention-rollout, as $E$ in Equation 3 will more strongly limit the model’s ability to condition on features other than the rationale tokens. We show a visual interpretation of this difference in Figure 5.
(a) Local
(b) Global
Figure 5: Illustration of local versus global attribution techniques as tools for ER. Local attributes to the input of the top-layer. Global attributes to the input tokens using the full model, and potentially constraining the model more strongly to follow human annotations.
Appendix B Plausibility Metrics
In order to measure how well the input attributions match the human annotated tokens we employ three metrics introduced in Fomicheva et al. (2021).
AUC Score.
Computes the Area Under the Receiver Operating Characteristic Curve score, which considers multiple threshold values for the attributions with human annotations as the target label. This metric has been used in past works DeYoung et al. (2020); Mathew et al. (2021); Fernandes et al. (2022); Resck et al. (2024).
Average Precision.
As mentioned in Fomicheva et al. (2021), AUC scores might be optimistic when dealing with unbalanced data. That is relevant when working with rationales, where the annotated tokens might correspond to a small portion of the input. Thus, we report the average precision score, which summarizes the average-precision curve as a weighted average of the precisions at the different thresholds,
$\textsc{AP}=\sum_{n}(R_{n}-R{n-1})~{}P_{n},$ (6)
where $P_{n}$ and $R_{n}$ correspond to the precision and recall values at a given threshold.
Recall@k.
Measures the average recall for a specific number of tokens $k$. It is calculated as
$r@k=\frac{|\{x\in r_{k}(x)~{}:~{}x<k\}|}{N},$ (7)
where $N$ is the number of annotated tokens for $x$, and $r(.)$ is a function that retrieves the rank of all annotated tokens.
Appendix C AUC Scores per Layer
As a complement to the results in Sections 6.2-6.4, Figure 6 shows average plausibility AUC scores across layers for baseline, ER with attention as the guided attribution technique (ER+Att and ER-C+Att), and ER with attention-rollout as the guided attribution technique (ER+AttR and ER-C+AttR). We can observe that the increase in plausibility for attention at the top-layer for ER+Att and ER-C+Att is not representative of changes in the plausibility of the model attributions. This observation holds for the attention attributions, and also for the ones obtained with ALTI and DecompX. On the other hand, attention-rollout (in the constrained setting) leads to changes in attributions for all layers, adding further evidence of its ability to impact the ER model to rely on plausible tokens.
Figure 6: Average AUC plausibility scores per-layer ($\uparrow$) computed with Attention, ALTI, and DecompX.
Appendix D Predicting OOD Performance
To complement the results discussed in Section 6.6, we obtain results on a wider range of OOD datasets and plausibility metrics criteria, and include also joint ER+Att and ER+AttR results. Results are shown in Figure 7. Similarly to Section 6.6, no in-domain criteria exhibits correlation with OOD classification performance.
(a) SST-Dev Classification as Predictor
(b) SST-Dev Plausibility as Predictor - AUC with Attention
(c) SST-Dev Plausibility as Predictor - AUC with Attention-Rollout
(d) SST-Dev Plausibility as Predictor - AUC with DecompX-Classifier
(e) SST-Dev Plausibility as Predictor - Average Precision with DecompX-Classifier
(f) SST-Dev Plausibility as Predictor - Recall@k with DecompX-Classifier
Figure 7: Scatter plot showing the relationship between in-domain (SST-Dev) measurements (F1-Macro classification scores and plausibility scores computed with different metrics and attribution techniques) for multiple datasets.
Appendix E Extra OOD Annotations
In order to further support our claims with respect to increased robustness to OOD conditions, we annotate 50 examples of the Yelp dataset. We first filter out examples that will result in more than 150 subtokens when using the tokenizer from the pre-trained encoder model we use during training (so that we can compute attributions with all available metrics, in particular ALTI (Ferrando et al., 2022) and DecompX (Modarressi et al., 2023), which we find to result in prohibitive runtime or OOM errors with long sequences). Then, we sample 55 reviews and label pairs at random from those, tokenize these examples with the English tokenizer from spaCy (Honnibal et al., 2020), and manually annotate the first five, selecting all sequences of text that offer evidence for the gold label. Those are use as few-shot examples for a prompt to Meta-Llama-3-8B-Instruct [^8] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct (AI@Meta, 2024) that outputs automatic annotations for the remaining 50 examples. Finally, we post-edit the automatic annotations whenever necessary.
Average model classification performance for this split of data is shown in Table 5.
Average F1-Macro
Baseline 93.71 $\pm$ 1.62
ER+Att 93.98 $\pm$ 1.80
ER+AttR 94.24 $\pm$ 2.07
ER-C+Att 94.64 $\pm$ 2.41
ER-C+AttR 92.64 $\pm$ 2.61
Table 5: F1-Macro scores ($\uparrow$) for the Yelp-50 split used to compute OOD attributions.
Appendix F Hyperparameters and Training
Hyperparameters.
We choose the combination of learning rate and maximum number of training epochs (which is relevant since we use a linear scheduler with 10% of training steps as warm-up) with the lowest cross-entropy loss over three runs with different seeds. For the explanation regularised approach we explore the same hyperparameters, plus the $\lambda$ weight. We select based on cross-entropy loss so that different values of $\lambda$ can be directly compared.
For the constrained approach, we first train a model (3 seeds) using $\mathcal{L}=\mathcal{L}_{\mathrm{expl}}$ and use the (rounded) average of the obtained minimum explanation train and validation losses as our bounds $b_{\mathrm{train}}$ and $b_{\mathrm{val}}$. Then, we choose the combination of learning rate and constrained optimizer learning rate that minimizes the average cross entropy loss.
Training.
During training we choose a model checkpoint based on average validation loss for the baseline and the joint explanation regularisation approach. This follows Joshi et al. (2022) and ensures that the explanation loss of the ER approach directly influences model selection.
For constrained optimization, we choose the model checkpoint with the lowest validation cross entropy loss, provided that the validation explanation loss is $\mathcal{L}_{\mathrm{expl}_{\mathrm{val}}}<1.1\times b_{\mathrm{val}}$. This ensures that we compare model checkpoints where the guided attribution technique is learning to predict the annotated rationales, according to the defined bound. Selecting a checkpoint based on the same criteria would have not been possible for the joint approach, as none of the runs converges to an explanation loss value that meets the defined bound. This further highlights the benefits of constrained optimization as analysis tool for our use case.
Unless mentioned otherwise, we report results over 15 seeds. All our experiments are developed using a single Nvidia A100 40GB GPU, and implemented with PyTorch (Paszke et al., 2019) and Pytorch Lightning (Falcon and The PyTorch Lightning team, 2019).
F.1 Hyperparameters
For the baseline we explore learning rate $\in\{2,3,5\times 10^{-5}\}$ and maximum number of epochs $\in\{15,25\}$. For the joint approach we explore the same space as the baseline model and $\lambda\in\{0.6,1.0,1.4\}$, following a set of choices aligned with previous work on ER for OOD robustness. For the constrained approach we explore learning rate $\in\{2,3,5\times 10^{-5}\}$ and constrained learning rate $\in\{1\times 10^{-1},5\times 10^{-2}\}$.
Final choices can be seen in Table 6. The trained models for Section 6.5 all use the same hyperparameters as the baseline model.
Shared Value
Batch Size 32
Optimizer AdamW
Weight Decay 0
$\beta$ (0.9, 0.98)
Scheduler Linear
Scheduler Warmup Steps 10%
Baseline Value
Learning Rate $3\times 10^{-5}$
Train Epochs 25
Joint Attention Value
Learning Rate $2\times 10^{-5}$
Train Epochs 25
$\mathbf{\lambda}$ 1.0
Joint Rollout Value
Learning Rate $3\times 10^{-5}$
Train Epochs 15
$\mathbf{\lambda}$ 1.0
Constrained Attention Value
Learning Rate $2\times 10^{-5}$
Train Epochs 25
Constrained Optimizer RMSprop
Constrained Learning Rate $5\times 10^{-2}$
$b_{\mathrm{train}}$ – Initial 0.035
$b_{\mathrm{train}}$ – Minimum 0.023
$b_{\mathrm{val}}$ 0.031
Constrained Rollout Value
Learning Rate $3\times 10^{-5}$
Train Epochs 25
Constrained Optimizer RMSprop
Constrained Learning Rate $1\times 10^{-1}$
$b_{\mathrm{train}}$ – Initial 0.03
$b_{\mathrm{train}}$ – Minimum 0.022
$b_{\mathrm{val}}$ 0.031
Table 6: Hyperparameters. (Loshchilov and Hutter, 2018)
F.2 Data
For in-domain and training data we use SST-2 (Socher et al., 2013), following the heuristic algorithm proposed in Carton et al. (2020) to obtain instance-level rationales [^9] https://github.com/BoulderDS/evaluating-human-rationales/blob/master/scripts/download_and_process_sst.py. For out-of-domain we use: Amazon-Reviews [^10] https://huggingface.co/datasets/McAuley-Lab/Amazon-Reviews-2023 (Hou et al., 2024), IMBD [^11] https://huggingface.co/datasets/stanfordnlp/imdb (Maas et al., 2011), Movies [^12] https://huggingface.co/datasets/eraser-benchmark/movie_rationales (Zaidan and Eisner, 2008; DeYoung et al., 2020), and Yelp [^13] https://huggingface.co/datasets/fancyzhx/yelp_polarity (Zhang et al., 2015). For some of the OOD evaluation data, the full test splits contain several thousand examples. In those cases we sample 5,000 examples. More details on the final data are described in Table 7. All data is in English.
Train Dev Test % Positive Avg. / Max Length Source Rationales
SST (Socher et al., 2013) 6,920 872 1,821 $\approx$ 50 % $\approx 25$ / $\approx 65$ Movies
Movies (Zaidan and Eisner, 2008; DeYoung et al., 2020) - - 399 50.0 % 786 / 1024 Movies
IMDB (Maas et al., 2011) - - 5,000 50.7 % 288 / 1024 Movies
$\times$
Amazon (Movies and TV) (Hou et al., 2024) - - 5,000 77.5 % 107 / 1024 Products
$\times$
Yelp (Zhang et al., 2015) - - 5,000 50.6 % 168 / 1024 Businesses
$\times$
Yelp-50∗∗ - - 50 50.0 % 84 / 151 Businesses
Table 7: Sentiment analysis datasets, including percentage of examples labeled as positive, average and maximum length of inputs (after truncating), source of the data, and whether rationales are provided. – This includes both dev and test examples. ∗∗ – This corresponds to the subset of annotated data described in Section E.
F.3 Guided Attribution Technique
Following Joshi et al. (2022), we scale the attributions $E(C_{\theta},x)$ by 100 when computing the explanation loss (Equation 3) and re-normalize them with softmax. We take it as part of the approach, and do so for both approaches that use attention as the guided attribution technique, ER+Att and ER-C+Att.
F.4 Evaluation Attribution Techniques
We keep all default choices part of the original work for DecompX, DecompX-C, and ALTI. For post-hoc attributions techniques that output ‘signed’ attribution scores in Section 5.1, i.e., IxG and DX-C, we take the absolute value of the attribution score.
F.5 Classifier
Following Joshi et al. (2022), classification is performed with a pre-trained Transformer encoder model, google/bigbird-roberta-base [^14] https://huggingface.co/google/bigbird-roberta-base (Zaheer et al., 2020). This is followed by a linear layer that uses as input the top-layer representation of the [CLS] token, with tanh as the non-linearity, and no classifier dropout.
AUC AP R@k
Att
AttR
IxG
ALTI
DX-C
DX
Att
AttR
IxG
ALTI
DX-C
DX
Att
AttR
IxG
ALTI
DX-C
DX
SST-Dev (ID)
Baseline 35.6 $\pm$ 2.9 27.9 $\pm$ 1.1 50.7 $\pm$ 1.5 59.0 $\pm$ 2.8 58.5 $\pm$ 1.6 57.5 $\pm$ 1.7 39.8 $\pm$ 2.0 35.6 $\pm$ 0.4 50.7 $\pm$ 1.0 61.2 $\pm$ 2.2 59.8 $\pm$ 1.6 58.3 $\pm$ 1.7 32.0 $\pm$ 1.9 28.0 $\pm$ 0.6 42.7 $\pm$ 0.9 52.3 $\pm$ 2.2 51.4 $\pm$ 1.3 50.4 $\pm$ 1.4
ER + Att 78.6 $\pm$ 2.8 27.9 $\pm$ 0.8 50.6 $\pm$ 1.8 64.5 $\pm$ 2.4 58.4 $\pm$ 0.9 57.8 $\pm$ 1.6 76.7 $\pm$ 2.2 35.7 $\pm$ 0.4 50.5 $\pm$ 1.5 64.1 $\pm$ 1.3 59.5 $\pm$ 0.9 58.0 $\pm$ 1.4 67.9 $\pm$ 2.3 28.3 $\pm$ 0.5 42.5 $\pm$ 1.2 55.5 $\pm$ 1.4 51.2 $\pm$ 0.9 50.3 $\pm$ 1.3
ER + AttR 37.0 $\pm$ 4.7 28.7 $\pm$ 0.6 50.8 $\pm$ 2.1 60.2 $\pm$ 3.0 58.6 $\pm$ 1.4 57.4 $\pm$ 2.0 40.8 $\pm$ 3.1 35.9 $\pm$ 0.3 50.5 $\pm$ 1.9 61.9 $\pm$ 2.1 60.2 $\pm$ 1.4 58.6 $\pm$ 2.0 32.9 $\pm$ 3.6 28.4 $\pm$ 0.4 42.8 $\pm$ 1.4 53.3 $\pm$ 2.1 51.8 $\pm$ 1.2 50.8 $\pm$ 1.6
ER-C + Att 87.6 $\pm$ 0.7 28.2 $\pm$ 0.5 51.5 $\pm$ 1.4 68.3 $\pm$ 3.6 57.6 $\pm$ 2.0 56.8 $\pm$ 1.5 84.8 $\pm$ 0.5 35.7 $\pm$ 0.2 51.2 $\pm$ 1.2 65.8 $\pm$ 2.1 58.6 $\pm$ 2.5 56.6 $\pm$ 1.6 76.3 $\pm$ 0.6 28.3 $\pm$ 0.4 43.2 $\pm$ 1.0 57.9 $\pm$ 2.2 50.6 $\pm$ 2.1 49.2 $\pm$ 1.4
ER-C + AttR 63.2 $\pm$ 8.4 89.1 $\pm$ 0.3 67.6 $\pm$ 1.5 87.9 $\pm$ 0.6 76.9 $\pm$ 0.9 81.3 $\pm$ 0.6 63.9 $\pm$ 7.7 85.6 $\pm$ 0.5 61.8 $\pm$ 1.1 84.9 $\pm$ 0.6 73.7 $\pm$ 0.7 77.0 $\pm$ 0.5 55.1 $\pm$ 7.6 77.1 $\pm$ 0.6 53.8 $\pm$ 1.1 76.5 $\pm$ 0.7 64.5 $\pm$ 0.8 67.9 $\pm$ 0.6
$\mathcal{L}_{\text{expl}}$ (A) 89.0 $\pm$ 0.4 29.6 $\pm$ 0.5 53.2 $\pm$ 1.6 78.2 $\pm$ 1.2 52.0 $\pm$ 0.7 53.0 $\pm$ 0.8 87.4 $\pm$ 0.3 36.0 $\pm$ 0.1 51.8 $\pm$ 0.9 73.8 $\pm$ 0.2 50.9 $\pm$ 1.0 50.0 $\pm$ 1.6 79.2 $\pm$ 0.3 28.6 $\pm$ 0.0 43.6 $\pm$ 0.6 65.0 $\pm$ 0.2 44.0 $\pm$ 0.5 43.6 $\pm$ 1.8
$\mathcal{L}_{\text{expl}}$ (R) 81.1 $\pm$ 2.4 89.4 $\pm$ 0.2 72.9 $\pm$ 1.4 88.2 $\pm$ 0.5 77.1 $\pm$ 2.5 83.3 $\pm$ 1.6 76.3 $\pm$ 1.8 86.7 $\pm$ 0.4 65.4 $\pm$ 0.6 84.0 $\pm$ 0.8 72.3 $\pm$ 1.6 77.6 $\pm$ 1.4 69.6 $\pm$ 2.1 78.6 $\pm$ 0.4 57.5 $\pm$ 0.6 76.3 $\pm$ 0.9 63.5 $\pm$ 1.6 69.4 $\pm$ 1.7
Movies (OOD)
Baseline 70.0 $\pm$ 2.5 49.6 $\pm$ 1.2 58.1 $\pm$ 1.4 - - - 27.7 $\pm$ 2.2 18.5 $\pm$ 0.3 22.6 $\pm$ 0.8 - - - 28.2 $\pm$ 2.7 16.6 $\pm$ 0.4 22.8 $\pm$ 1.0 - - -
ER + Att 74.6 $\pm$ 2.0 49.1 $\pm$ 0.7 58.1 $\pm$ 0.7 - - - 34.9 $\pm$ 2.9 18.4 $\pm$ 0.2 22.7 $\pm$ 0.5 - - - 35.3 $\pm$ 2.9 16.5 $\pm$ 0.2 23.1 $\pm$ 0.6 - - -
ER + AttR 70.1 $\pm$ 1.7 50.2 $\pm$ 1.1 58.0 $\pm$ 1.0 - - - 27.5 $\pm$ 1.3 18.7 $\pm$ 0.3 22.5 $\pm$ 0.5 - - - 28.0 $\pm$ 1.5 16.8 $\pm$ 0.4 22.7 $\pm$ 0.6 - - -
ER-C + Att 68.8 $\pm$ 1.6 48.1 $\pm$ 0.6 56.0 $\pm$ 1.2 - - - 31.7 $\pm$ 1.6 18.2 $\pm$ 0.1 21.5 $\pm$ 0.6 - - - 32.6 $\pm$ 1.4 16.3 $\pm$ 0.2 21.6 $\pm$ 0.8 - - -
ER-C + AttR 71.1 $\pm$ 2.0 62.8 $\pm$ 1.3 59.3 $\pm$ 1.1 - - - 31.3 $\pm$ 2.4 26.1 $\pm$ 0.8 23.1 $\pm$ 0.7 - - - 32.1 $\pm$ 2.7 26.6 $\pm$ 1.0 23.6 $\pm$ 0.9 - - -
$\mathcal{L}_{\text{expl}}$ (A) 61.0 $\pm$ 2.4 48.2 $\pm$ 1.5 48.0 $\pm$ 0.3 - - - 25.6 $\pm$ 0.6 18.0 $\pm$ 0.3 18.3 $\pm$ 0.1 - - - 27.4 $\pm$ 0.3 16.1 $\pm$ 0.4 17.0 $\pm$ 0.0 - - -
$\mathcal{L}_{\text{expl}}$ (R) 59.4 $\pm$ 1.3 62.5 $\pm$ 0.4 55.2 $\pm$ 0.8 - - - 23.8 $\pm$ 0.3 25.8 $\pm$ 0.2 21.7 $\pm$ 0.4 - - - 22.7 $\pm$ 0.8 26.6 $\pm$ 0.2 21.2 $\pm$ 0.5 - - -
Yelp-50 (OOD)
Baseline 51.9 $\pm$ 2.5 37.1 $\pm$ 1.1 51.5 $\pm$ 1.4 66.2 $\pm$ 2.1 59.5 $\pm$ 2.0 60.3 $\pm$ 1.4 28.5 $\pm$ 2.2 21.8 $\pm$ 0.3 31.1 $\pm$ 1.2 48.4 $\pm$ 3.9 41.7 $\pm$ 2.5 39.5 $\pm$ 2.1 23.2 $\pm$ 3.0 14.3 $\pm$ 0.6 26.3 $\pm$ 1.4 41.7 $\pm$ 2.7 36.6 $\pm$ 2.4 36.3 $\pm$ 1.7
ER + Att 69.7 $\pm$ 2.0 36.4 $\pm$ 1.1 50.7 $\pm$ 1.2 67.4 $\pm$ 1.9 59.1 $\pm$ 0.9 59.3 $\pm$ 1.5 53.1 $\pm$ 1.8 21.7 $\pm$ 0.3 30.8 $\pm$ 1.2 50.5 $\pm$ 2.2 40.6 $\pm$ 1.5 38.0 $\pm$ 1.7 46.8 $\pm$ 2.2 14.0 $\pm$ 0.5 25.6 $\pm$ 1.4 43.9 $\pm$ 2.1 35.7 $\pm$ 1.4 35.2 $\pm$ 1.7
ER + AttR 52.5 $\pm$ 3.0 37.6 $\pm$ 1.2 51.7 $\pm$ 0.8 67.1 $\pm$ 1.5 59.8 $\pm$ 1.2 60.5 $\pm$ 1.3 28.9 $\pm$ 2.7 22.0 $\pm$ 0.3 31.4 $\pm$ 1.2 49.8 $\pm$ 2.8 42.6 $\pm$ 1.7 40.5 $\pm$ 2.1 23.5 $\pm$ 3.9 14.4 $\pm$ 0.7 26.6 $\pm$ 1.3 42.8 $\pm$ 2.1 37.0 $\pm$ 1.3 36.7 $\pm$ 1.7
ER-C + Att 76.0 $\pm$ 1.2 37.4 $\pm$ 1.2 51.3 $\pm$ 1.4 69.4 $\pm$ 2.7 58.6 $\pm$ 2.9 58.9 $\pm$ 2.5 56.4 $\pm$ 1.5 21.9 $\pm$ 0.3 31.1 $\pm$ 1.2 49.5 $\pm$ 2.7 39.2 $\pm$ 3.4 36.8 $\pm$ 2.8 50.9 $\pm$ 1.6 14.4 $\pm$ 0.7 26.2 $\pm$ 1.6 43.6 $\pm$ 2.8 35.4 $\pm$ 3.2 34.5 $\pm$ 2.9
ER-C + AttR 67.2 $\pm$ 5.8 69.4 $\pm$ 1.5 58.8 $\pm$ 1.4 76.7 $\pm$ 1.6 65.5 $\pm$ 1.5 67.9 $\pm$ 1.6 49.4 $\pm$ 7.5 49.5 $\pm$ 1.5 37.8 $\pm$ 1.6 58.3 $\pm$ 1.8 47.7 $\pm$ 1.3 49.0 $\pm$ 1.3 43.4 $\pm$ 7.1 43.8 $\pm$ 1.4 34.1 $\pm$ 1.6 51.8 $\pm$ 1.8 42.3 $\pm$ 1.5 43.4 $\pm$ 1.5
$\mathcal{L}_{\text{expl}}$ (A) 70.1 $\pm$ 2.1 38.5 $\pm$ 1.9 47.1 $\pm$ 0.7 62.5 $\pm$ 1.3 51.3 $\pm$ 1.3 49.4 $\pm$ 3.1 50.3 $\pm$ 2.5 21.8 $\pm$ 0.7 27.9 $\pm$ 0.4 41.2 $\pm$ 1.1 30.0 $\pm$ 1.0 27.8 $\pm$ 1.6 45.4 $\pm$ 1.7 14.6 $\pm$ 1.1 23.6 $\pm$ 0.4 36.3 $\pm$ 1.3 25.9 $\pm$ 1.0 23.1 $\pm$ 2.3
$\mathcal{L}_{\text{expl}}$ (R) 62.6 $\pm$ 2.9 68.6 $\pm$ 1.0 56.8 $\pm$ 1.7 66.5 $\pm$ 1.3 59.0 $\pm$ 1.7 61.6 $\pm$ 0.2 39.3 $\pm$ 2.9 47.1 $\pm$ 1.8 35.2 $\pm$ 1.4 43.8 $\pm$ 1.7 37.1 $\pm$ 1.3 39.0 $\pm$ 0.4 35.7 $\pm$ 4.0 42.1 $\pm$ 2.0 31.9 $\pm$ 1.2 39.5 $\pm$ 2.1 33.7 $\pm$ 1.4 35.6 $\pm$ 0.8
Table 8: Attributions to Annotated Tokens. AUC, Average Precision and Recall@k Scores ($\uparrow$) for SST-Dev (ID), Movies (OOD), and Yelp-50 (OOD) datasets, for the Baseline (Baseline), ER model guided with Attention (ER + Att), ER model guided with Attention-Rollout (ER + AttR), and the corresponding constrained counterparts (ER-C + X). Attributions are computed using Attention (Att), Attention + Rollout (AttR), InputxGradient (IxG), ALTI Aggregated (ALTI), and DecompX (with) Classifier (DX and DX-C) as the attribution techniques. For Movies, we report only the techniques that can be applied to longer sequences.